# Revisiting Applicable and Comprehensive Knowledge Tracing in Large-Scale Data

### v1

@@ 3.3.3 IRT Prediction & Decompose
Table 1: One-step prediction performance of DKT2 and 17 baselines from different categories. The best result is in bold, the second best is underlined. ✓ indicates strong applicability and comprehensiveness, ✗ indicates weak applicability and comprehensiveness, ✝ indicates strong applicability but weak comprehensiveness. - indicates the model fails to be applied to such a large-scale dataset.
3.3.4 Integrated Knowledge Fusion
Finally, DKT2 predicts the student’s comprehensive knowledge states $K\!S_{2:t+1}$:
$\begin{split}K\!S_{2:t+1}=\sigma(\text{ReLU}(X_{2:t+1}W_{1}+b_{1})W_{2}+b_{2}),\end{split}$ (7)
Eq. 7 can be further represented as:
$\begin{split}K\!S_{i}=(\hat{r}_{i}^{1},\hat{r}_{i}^{2},\dots,\hat{r}_{i}^{n}),2\leq i\leq t+1,\end{split}$ (8)
where $\hat{r}_{i}^{j}$ represents the prediction score of DKT2 for concept $j$ at time step $i$. The comprehensive output of DKT2 enables the prediction of multiple concepts at the same time step, whereas models like AKT can only predict $\hat{r}_{i}^{c_{i}}$ at time step $i$. We will analyze the multi-concept prediction scenario in detail in Section 4.3.2, where some unexpected results have been discovered.
3.3.5 Model Training
$\mathcal{L}_{\text{bce}}=-\sum_{i=2}^{t+1}{r_{i}\text{log}({\hat{r}}_{i})+(1-r_{i})\text{log}(1-{\hat{r}}_{i})}.$ (9)
4 Experiments
Our goal is to answer the following research questions:
  • RQ1: How does DKT2 perform compared to 17 baselines from 8 different categories under applicable input and comprehensive output settings?
@@ 4 Experiments
Table 2: Multi-step prediction performance of DKT2 and several representative baselines on Assist17.
4.1 Experimental Setup
4.1.1 Datasets
We conduct extensive experiments on three of the latest large-scale benchmark datasets from different platforms: Assist17 Feng et al. (2009), EdNet Choi et al. (2020), and Comp Hu et al. (2023). Details of the datasets are provided in Appendix A.4.
4.1.2 Baselines
To comprehensively and systematically evaluate the performance of DKT2 and analyze the impact of input-output settings on KT models, we compare DKT2 with 17 DLKT baselines from 8 categories, as mentioned in Section 2. Detailed descriptions of the aforementioned DLKT baselines can be found in Appendix A.5.
4.1.3 Implementation
Similar to CL4KT Lee et al. (2022), we employ five-fold cross-validation, with folds divided by students. 10% of the training set is used for model evaluation and also for the early stopping strategy: if the AUC does not improve within 10 epochs during the 300 epochs, the training will be stopped. The averages across five test folds are reported. We focus on the most recent 100 interactions (history length) for each student, as this latest information is crucial for future predictions. During training, all models are trained using the Adam optimizer Kingma and Ba (2017) with the following unified settings: batch size is fixed at 512, learning rate is 0.001, dropout rate is 0.05, and embedding dimension is 64. The seed is set to 12405 to reproduce experimental results. Similar to existing DLKT research, our evaluation metrics include two classification metrics, Area Under the ROC Curve (AUC) and Accuracy (ACC), and one regression metric, Root Mean Square Error (RMSE).
@@ 4.1.3 Implementation
Figure 3: The prediction performance of DKT2 and several representative baselines on Assist17 with different history lengths.
4.2 Applicable and Comprehensive Performance Comparison (RQ1)
Under applicable input and comprehensive output settings [^3] See Appendix A.6 for input and output settings details., we evaluate common prediction tasks in KT Liu et al. (2022): 1) one-step prediction, 2) multi-step prediction, and 3) prediction with varying history lengths.
4.2.1 One-step Prediction
KT’s one-step prediction can provide immediate feedback for ITS and be used for short-term adjustments of personalized learning paths Corbett and Anderson (1994). Table 1 shows the one-step prediction performance of DKT2 and 17 baselines from 8 different categories in three large-scale datasets. Overall, in this fair large-scale data competition, our DKT2 has emerged as the final winner by a narrow margin. We observe:
  • Compared to previous research Ghosh et al. (2020), under the input-output settings, attention-based models like AKT still generally outperform deep sequential models like DKT, suggesting that attention-based models like AKT may be less affected by these settings.
  • The recently proposed Mamba4KT performs well on Assist17, but underperforms compared to DKT on larger-scale datasets like EdNet and Comp. This may be due to mamba’s poorer performance in context learning in large-scale experiments, which is consistent with previous research findings Waleffe et al. (2024).
  • Our proposed DKT2 performs almost the best across all metrics. This performance improvement can be attributed to the superiority of DKT2, which includes the exponential activation function in sLSTM that helps improve memory and forgetting processes, and the matrix memory introduced in mLSTM that gives DKT2 advantages in large-scale applications and long sequence processing.
@@ 4.2.1 One-step Prediction
Setting Metric AKT simpleKT FoLiBiKT sparseKT DTransformer DKVMN CL4KT Deep-IRT
@@
4.2.2 Multi-step Prediction
KT’s accurate multi-step prediction not only provides valuable feedback for selecting and constructing personalized learning materials, but also assists ITS in flexibly adjusting future curriculum based on student needs Liu et al. (2023b). Table 2 and Table 6 in Appendix A.7.1 show the multi-step (step=5, 10, 15, 20) prediction performance of DKT2 and several representative baselines from different categories. The main observations are as follows: (1) As the prediction steps increase, the performance of all models consistently decreases. This is due to error accumulation, meaning that small errors in one-step prediction can accumulate over multiple steps, leading to a decrease in multi-step prediction performance. (2) Compared to one-step prediction, attention-based models perform well in multi-step prediction. This is because the attention mechanism can capture long-distance dependencies, making its advantages more apparent. In contrast, Mamba4KT performs poorly, as mamba-based models are highly dependent on context Lieber et al. (2024) and are more susceptible to error accumulation. (3) Our DKT2 generally outperforms all models in multi-step prediction. We can similarly attribute this to the exponential activation function introduced in sLSTM of DKT2, which can mitigate error accumulation by modifying storage decisions, as it allows the model to update its internal state at each step, while the matrix memory introduced in mLSTM provides support for large-capacity storage space.
4.2.3 Varying-history-length Prediction
Analyzing the impact of different history lengths can help ITS better understand students’ knowledge acquisition and forgetting processes, thereby improving teaching strategies. Figure 3 and Figures 6 and 7 in Appendix A.7.2 show the prediction performance of DKT2 and several representative baselines with different history lengths. From these, we can observe: 1) As the history length increases, the prediction performance of almost all models generally improves, as longer sequences provide more historical information. Surprisingly, Mamba4KT’s performance consistently decreases. A possible reason is that mamba-based models are better at capturing local temporal dependencies but may struggle to effectively capture long-distance dependencies within longer sequences. 2) Notably, DKT, using only one LSTM, can maintain a strong ranking position across different history lengths, further encouraging KT researchers to design simple yet effective models Liu et al. (2023b). 3) Our DKT2 maintains optimal performance across different history lengths, with more significant performance improvements as the history length increases. This is not only due to the increased storage capacity of mLSTM but also related to sLSTM providing a broader output range as the sequence length increases.
4.3 In-Depth Analysis (RQ2 & RQ3)
@@ 4.3.2 Multi-concept Prediction
Due to space limitation, the discussion on multi-concept prediction is in Appendix A.7.4, where we have discovered an interesting phenomenon.
@@ 4.3.3 Ablation Study
Figure 4 and Figure 9 in Appendix A.7.5 illustrate the impact of different components on DKT2. “w/o. Rasch" indicates the removal of Rasch embedding from DKT2 (setting $d_{q_{t}}$ to 00 in Eq. 2), “w/o. IRT" represents the removal of the IRT module, “w/o. IKF" means DKT2 ignores the integrated knowledge fusion, while “w/o. sLSTM" and “w/o. mLSTM" denote the removal of sLSTM block and mLSTM block, respectively. The results show that DKT2 achieves the highest AUC scores across all datasets compared to other variants, demonstrating the importance of each component on DKT2. Notably, “w/o. mLSTM" generally outperforms DKT2 on ACC and RMSE scores on Assist17, which is due to mLSTM’s inability to demonstrate significant advantages in small-scale data, as evidenced by its poorer performance on larger datasets, EdNet and Comp.
@@ 5 Conclusion
This paper introduces DKT2, an applicable and comprehensive DLKT model that addresses key limitations of deep sequential models like DKT. By leveraging xLSTM, the Rasch model, and Item Response Theory (IRT), DKT2 effectively balances predictive performance with practical applicability.
@@ 6 Limitations
Appendix A Appendix
@@
  • Summary of Strengths and Weaknesses of DLKT Models from 2015-2024 in Terms of Applicability and Comprehensiveness (A.1)
  • Detailed Introduction to LSTM (A.2)
  • Detailed Introduction to xLSTM (A.3)
  • Dataset Description and Processing Methods (A.4)
  • Baseline Description (A.5)
  • Input and Output Settings (A.6)
  • Additional Experimental Results (A.7)
A.1 Summary of Strengths and Weaknesses of DLKT Models from 2015-2024 in Terms of Applicability and Comprehensiveness
@@
Table 4 summarizes the strengths and weaknesses in terms of applicability and comprehensiveness of DLKT models in top AI/ML conferences/journals from 2015-2024.
Conference/Journal Model Applicability Comprehensiveness
AAAI KTM Vie and Kashima (2019)
QIKT Chen et al. (2023)
CIKM RKT Pandey and Srivastava (2020)
FoLiBiKT Im et al. (2023)
ICDM DKT-DSC Minn et al. (2018)
SKT Tong et al. (2020)
CAKT Yang et al. (2022)
ICLR simpleKT Liu et al. (2023b)
KDD AKT Ghosh et al. (2020)
LPKT Shen et al. (2021)
LBKT Xu et al. (2023)
MM ATKT Guo et al. (2021)
ABQR Sun et al. (2023)
NIPS DKT Piech et al. (2015)
PKDD GIKT Yang et al. (2021)
GMKT Zhao and Sahebi (2023)
SIGIR SKVMN Abdelrahman and Wang (2019)
CKT Shen et al. (2020)
IEKT Long et al. (2021)
DIMKT Shen et al. (2022b)
sparseKT Huang et al. (2023)
TKDE LPKT-S Shen et al. (2022a)
WSDM HawkesKT Wang et al. (2021)
CoKT Long et al. (2022)
WWW DKVMN Zhang et al. (2017)
DKT-F Nagatani et al. (2019)
CL4KT Lee et al. (2022)
DTransformer Yin et al. (2023)
AT-DKT Liu et al. (2023a)
MIKT Sun et al. (2024)
Table 4: Summary of the applicability and comprehensiveness of DLKT models in top AI/ML conferences/journals from 2015-2024. ✔ and ✘ indicate strong and weak applicability and comprehensiveness, respectively.
A.2 Detailed Introduction to LSTM
@@
Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber (1997) overcomes the short-term memory limitations of Recurrent Neural Networks (RNN) Elman (1990) caused by the vanishing gradient [^4] The vanishing gradient refers to the phenomenon where, during model training, as time step increases, the gradient is continuously multiplied by the weight matrix during backpropagation, potentially causing it to shrink rapidly towards zero, resulting in very slow weight updates in the network. Hochreiter (1991); Hochreiter et al. (2001) by introducing cell state and gating mechanisms into the network. Figure 5 shows the architecture of LSTM at time step $t$. The core concepts of LSTM include cell state and various gate structures. The cell state acts as a pathway for transmitting relevant information, allowing information to be passed along the sequence chain, which can be viewed as the network’s memory. Theoretically, during sequence processing, the cell state can continuously carry relevant information. Thus, information obtained at earlier time steps can be transmitted to cells at later time steps, which helps mitigate the impact of short-term memory. Additionally, LSTM addresses the short-term memory issue of RNNs by introducing internal gating mechanisms (i.e., forget gate Gers et al. (2000), input gate, and output gate) to regulate information flow. Specifically, LSTM uses the Tanh activation function (with output values always in the range (-1, 1)) to help regulate the neural network output and employs the Sigmoid activation function in its gate structures. The Sigmoid function is similar to the Tanh function, but its output range is (0, 1), which aids in updating or forgetting data, as any number multiplied by 0 becomes 0 (this information is forgotten), and any number multiplied by 1 remains unchanged (this information is fully preserved). This allows the network to understand which data is unimportant and should be forgotten, and which data is important and should be preserved. The cell state update rule (i.e., the constant error carousel Hochreiter and Schmidhuber (1996)) for LSTM at time step $t$ is:
@@
In summary, the forget gate in LSTM determines which relevant information from previous time steps should be preserved, the input gate decides which important information from the current input should be added, and the output gate determines the next hidden state. Previous work Greff et al. (2016) has shown that each gate structure is crucial. Recently, LSTM has been revisited and greatly improved, with the revised LSTM known as xLSTM Beck et al. (2024). xLSTM enhances the traditional LSTM structure, aiming to improve LSTM’s performance and scalability with large-scale data. Subsequently, a series of studies on xLSTM have been applied to various fields such as computer vision Alkin et al. (2024); Zhu et al. (2024); Dutta et al. (2024) and time series Alharthi and Mahmood (2024).
Figure 5: Architecture of LSTM.
A.3 Detailed Introduction to xLSTM
A.3.1 Stabilized Long Short-Term Memory
To enable LSTM to revise storage decisions, sLSTM introduces an exponential activation function along with normalizer state and stabilization. Unlike the Sigmoid activation function (i.e., S-shaped function) mentioned in Appendix A.2, where it becomes very challenging for the model to decide what to forget or retain as input values get higher, sLSTM uses an exponential function instead, providing a broader output range, indicating that sLSTM can better revise storage decisions. However, after introducing the exponential function, output values tend to surge as input values increase and do not naturally normalize outputs between 0 and 1 as the Sigmoid function does. Therefore, sLSTM introduces normalizer state, which is a function of the forget gate and input gate, to normalize the hidden state. The update rule for the sLSTM cell state at time step $t$ is:
@@ A.3.1 Stabilized Long Short-Term Memory
Moreover, since the exponential activation function can easily cause overflow for large values, to prevent the exponential function from disrupting the forget gate and input gate, sLSTM uses an additional state $m_{t}$ Milakov and Gimelshein (2018), which appears in logarithmic form, to counteract the effect of the exponential function and introduce stability:
@@
A.3.2 Matrix Long Short-Term Memory
To enhance LSTM’s memory ability to capture more complex data relationships and patterns, mLSTM introduces a matrix $C\in\mathbb{R}^{d\times d}$ to replace the scalar cell state $c\in\mathbb{R}$. Additionally, since LSTM is designed to process sequential data, which means it needs to process the output of the previous input in the sequence to handle the current input, this hinders parallelization and is the main culprit leading to the Transformer era. Therefore, mLSTM abandons this design concept. Specifically, mLSTM adopts the setting of Bidirectional Associative Memories (BAMs) Kohonen (1972); Anderson et al. (1977): at time step $t$, mLSTM stores a pair of vectors, key $k_{t}\in\mathbb{R}^{d}$ and value $v_{t}\in\mathbb{R}^{d}$. At time step $t+\tau$, the value $v_{t}$ is retrieved through a query vector $q_{t+\tau}\in\mathbb{R}^{d}$. mLSTM uses a covariance update rule ($C_{t}=C_{t-1}+v_{t}\,k_{t}^{\top}$) to store the key-value pair. The covariance update rule is equivalent to the Fast Weight Programmer Schmidhuber (1992). Later, a new variant has emerged Ba et al. (2016a): a constant decay rate multiplied by $C_{t-1}$ and a constant learning rate multiplied by $v_{t}\,k_{t}^{\top}$. Similarly, in mLSTM, the forget gate corresponds to the decay rate, while the input gate corresponds to the learning rate. Furthermore, since the dot product between the query input and the normalizer state may approach zero, mLSTM uses the absolute value of the dot product and sets a lower bound to a threshold (e.g., 1). The cell state update rule for mLSTM is:

### v2

@@ 4.2 Applicable and Comprehensive Performance Comparison (RQ1)
Table 1: One-step prediction performance of DKT2 and 18 baselines from different categories. The best result is in bold, the second best is underlined. ✓ indicates strong applicability and comprehensiveness, ✗ indicates weak applicability and comprehensiveness, ✝ indicates strong applicability but weak comprehensiveness. - indicates the model fails to be applied to such a large-scale dataset, resulting in a program crash.
4.2.1 One-step Prediction
KT’s one-step prediction can provide immediate feedback for ITS and be used for short-term adjustments of personalized learning paths [16]. Table 1 shows the one-step prediction performance of DKT2 and 18 baselines from 8 different categories in three large-scale datasets. Overall, in this fair large-scale data competition, our DKT2 has emerged as the final winner by a narrow margin. We observe:
  • Compared to previous research [26], under the input-output settings, attention-based models like AKT still generally outperform deep sequential models like DKT, suggesting that attention-based models like AKT may be less affected by these settings.
  • The recently proposed Mamba4KT performs well on Assist17, but underperforms compared to DKT on larger-scale datasets like EdNet and Comp. This may be due to mamba’s poorer performance in context learning in large-scale experiments, which is consistent with previous research findings [84].
  • Our proposed DKT2 performs almost the best on all metrics across all datasets. This performance improvement can be attributed to the superiority of DKT2, which includes the exponential activation function in sLSTM that helps improve memory and forgetting processes, and the matrix memory introduced in mLSTM that gives DKT2 advantages in large-scale applications and long sequence processing.
@@ 4.2.1 One-step Prediction
Table 2: Multi-step prediction performance of DKT2 and several representative baselines on Assist17. The results for EdNet and Comp can be found in Appendix 0.A.6.1.
4.2.2 Multi-step Prediction
KT’s accurate multi-step prediction not only provides valuable feedback for selecting and constructing personalized learning materials, but also assists ITS in flexibly adjusting future curriculum based on student needs [54]. Table 2 and Table 7 in Appendix 0.A.6.1 show the multi-step (step=5, 10, 15, 20) prediction performance of DKT2 and several representative baselines from different categories. The main observations are as follows: (1) As the prediction steps increase, the performance of all models consistently decreases. This is due to error accumulation, meaning that small errors in one-step prediction can accumulate over multiple steps, leading to a decrease in multi-step prediction performance. (2) Compared to one-step prediction, attention-based models perform well in multi-step prediction. This is because the attention mechanism can capture long-distance dependencies, making its advantages more apparent. In contrast, Mamba4KT performs poorly, as mamba-based models are highly dependent on context [48] and are more susceptible to error accumulation. (3) Our DKT2 generally outperforms all models in multi-step prediction. We can similarly attribute this to the exponential activation function introduced in sLSTM of DKT2, which can mitigate error accumulation by modifying storage decisions, as it allows the model to update its internal state at each step, while the matrix memory introduced in mLSTM provides support for large-capacity storage space.
@@ 4.2.2 Multi-step Prediction
Figure 4: The prediction performance of DKT2 and several representative baselines on Assist17 with different history lengths. The results for EdNet and Comp are in Appendix 0.A.6.1.
4.2.3 Varying-history-length Prediction
Analyzing the impact of different history lengths can help ITS better understand students’ knowledge acquisition and forgetting processes, thereby improving teaching strategies. Fig. 4, Fig. 7 and Fig. 8 in Appendix 0.A.6.2 show the prediction performance of DKT2 and several representative baselines with different history lengths. From these, we can observe: 1) As the history length increases, the prediction performance of almost all models generally improves, as longer sequences provide more historical information. Surprisingly, Mamba4KT’s performance consistently decreases. A possible reason is that mamba-based models are better at capturing local temporal dependencies but may struggle to effectively capture long-distance dependencies within longer sequences. 2) Notably, DKT, using only one LSTM, can maintain a strong ranking position across different history lengths, further encouraging KT researchers to design simple yet effective models [54]. 3) Our DKT2 maintains optimal performance across different history lengths, with more significant performance improvements as the history length increases. This is not only due to the increased storage capacity of mLSTM but also related to sLSTM providing a broader output range as the sequence length increases.
4.3 In-Depth Analysis (RQ2 & RQ3)
@@
Setting Metric AKT simpleKT FoLiBiKT sparseKT DTransformer stableKT DKVMN CL4KT Deep-IRT
@@
[…]
@@ 4.3.1 Analysis of Different Input Settings
Dataset Assist17 EdNet Comp
[…]
DKT 0.5841 0.5787 0.4913 0.6600 0.6225 0.4775 0.7091 0.8045 0.3806
SAKT 0.5596 0.5534 0.5048 0.6546 0.6198 0.4788 0.6994 0.8037 0.3824
AKT 0.6185 0.6040 0.4862 0.6649 0.6241 0.4765 0.7054 0.8039 0.3815
DKVMN 0.5730 0.5701 0.4964 0.6572 0.6205 0.4781 0.7050 0.8034 0.3817
ATKT 0.6205 0.6077 0.4836 0.6639 0.6229 0.4768 0.7111 0.8049 0.3802
CL4KT 0.5892 0.5911 0.4890 - - - 0.7044 0.8032 0.3820
DKT2 0.6174 0.6041 0.4872 0.6646 0.6243 0.4768 0.7064 0.8047 0.3810
@@ 4.3.3 Ablation Study
Fig. 5 and Fig. 9 in Appendix 0.A.6.4 illustrate the impact of different components on DKT2. “w/o. Rasch" indicates the removal of Rasch embedding from DKT2 (setting $d_{q_{t}}$ to 00 in Eq. 2), “w/o. IRT" represents the removal of the IRT module, “w/o. IKF" means DKT2 ignores the integrated knowledge fusion, while “w/o. sLSTM" and “w/o. mLSTM" denote the removal of sLSTM block and mLSTM block, respectively. The results show that DKT2 achieves the highest AUC scores across all datasets compared to other variants, demonstrating the importance of each component on DKT2. Notably, “w/o. mLSTM" generally outperforms DKT2 on ACC and RMSE scores on Assist17, which is due to mLSTM’s inability to demonstrate significant advantages in small-scale data, as evidenced by its poorer performance on larger datasets, EdNet and Comp.
@@ 5 Conclusion
This paper introduces DKT2, an applicable and comprehensive DLKT model that addresses key limitations of deep sequential models like DKT. By leveraging xLSTM, the Rasch model, and Item Response Theory (IRT), DKT2 effectively balances predictive performance with practical applicability. Our extensive experiments across three large-scale datasets demonstrate DKT2’s superiority over 18 baseline models in various prediction tasks, highlighting its robustness and potential for real-world educational applications.
@@ 6 Limitations
Acknowledgments
This research was supported by grants from the "Pioneer" and "Leading Goose" R&D Program of Zhejiang (2025C02022), National Natural Science Foundation of China (No.62307032), and Shanghai Rising-Star Program (23QA1409000).
Appendix 0.A Appendix
@@
  • Summary of DLKT Models from 2015-2025 in Terms of Applicability and Comprehensiveness (0.A.1)
  • Detailed Introduction to LSTM (0.A.2)
  • Detailed Introduction to xLSTM (0.A.3)
  • Dataset Description and Processing Methods (0.A.4)
  • Baseline Description (0.A.5)
  • Additional Experimental Results (0.A.6)
0.A.1 Summary of DLKT Models from 2015-2025 in Terms of Applicability and Comprehensiveness
@@
Table 5 summarizes the DLKT models in terms of applicability and comprehensiveness in top AI/ML conferences/journals from 2015-2025.
0.A.2 Detailed Introduction to LSTM
Figure 6: Architecture of LSTM.
@@ 0.A.2 Detailed Introduction to LSTM
Long Short-Term Memory (LSTM) [37] overcomes the short-term memory limitations of Recurrent Neural Networks (RNN) [22] caused by the vanishing gradient [^3] The vanishing gradient refers to the phenomenon where, during model training, as time step increases, the gradient is continuously multiplied by the weight matrix during backpropagation, potentially causing it to shrink rapidly towards zero, resulting in very slow weight updates in the network. [34, 35] by introducing cell state and gating mechanisms into the network. Fig. 6 shows the architecture of LSTM at time step $t$. The core concepts of LSTM include cell state and various gate structures. The cell state acts as a pathway for transmitting relevant information, allowing information to be passed along the sequence chain, which can be viewed as the network’s memory. Theoretically, during sequence processing, the cell state can continuously carry relevant information. Thus, information obtained at earlier time steps can be transmitted to cells at later time steps, which helps mitigate the impact of short-term memory. Additionally, LSTM addresses the short-term memory issue of RNNs by introducing internal gating mechanisms (i.e., forget gate [25], input gate, and output gate) to regulate information flow. Specifically, LSTM uses the Tanh activation function (with output values always in the range (-1, 1)) to help regulate the neural network output and employs the Sigmoid activation function in its gate structures. The Sigmoid function is similar to the Tanh function, but its output range is (0, 1), which aids in updating or forgetting data, as any number multiplied by 0 becomes 0 (this information is forgotten), and any number multiplied by 1 remains unchanged (this information is fully preserved). This allows the network to understand which data is unimportant and should be forgotten, and which data is important and should be preserved. The cell state update rule (i.e., the constant error carousel [36]) for LSTM at time step $t$ is:
@@
In summary, the forget gate in LSTM determines which relevant information from previous time steps should be preserved, the input gate decides which important information from the current input should be added, and the output gate determines the next hidden state. Previous work [27] has shown that each gate structure is crucial. Recently, LSTM has been revisited and greatly improved, with the revised LSTM known as xLSTM [8]. xLSTM enhances the traditional LSTM structure, aiming to improve LSTM’s performance and scalability with large-scale data. Subsequently, a series of studies on xLSTM have been applied to various fields such as computer vision [4, 106, 21] and time series [3].
0.A.3 Detailed Introduction to xLSTM
0.A.3.1 Stabilized Long Short-Term Memory
To enable LSTM to revise storage decisions, sLSTM introduces an exponential activation function along with normalizer state and stabilization. Unlike the Sigmoid activation function (i.e., S-shaped function) mentioned in Appendix 0.A.2, where it becomes very challenging for the model to decide what to forget or retain as input values get higher, sLSTM uses an exponential function instead, providing a broader output range, indicating that sLSTM can better revise storage decisions. However, after introducing the exponential function, output values tend to surge as input values increase and do not naturally normalize outputs between 0 and 1 as the Sigmoid function does. Therefore, sLSTM introduces normalizer state, which is a function of the forget gate and input gate, to normalize the hidden state. The update rule for the sLSTM cell state at time step $t$ is:
@@ 0.A.3.1 Stabilized Long Short-Term Memory
Moreover, since the exponential activation function can easily cause overflow for large values, to prevent the exponential function from disrupting the forget gate and input gate, sLSTM uses an additional state $m_{t}$ [61], which appears in logarithmic form, to counteract the effect of the exponential function and introduce stability:
@@
0.A.3.2 Matrix Long Short-Term Memory
To enhance LSTM’s memory ability to capture more complex data relationships and patterns, mLSTM introduces a matrix $C\in\mathbb{R}^{d\times d}$ to replace the scalar cell state $c\in\mathbb{R}$. Additionally, since LSTM is designed to process sequential data, which means it needs to process the output of the previous input in the sequence to handle the current input, this hinders parallelization and is the main culprit leading to the Transformer era. Therefore, mLSTM abandons this design concept. Specifically, mLSTM adopts the setting of Bidirectional Associative Memories (BAMs) [44, 5]: at time step $t$, mLSTM stores a pair of vectors, key $k_{t}\in\mathbb{R}^{d}$ and value $v_{t}\in\mathbb{R}^{d}$. At time step $t+\tau$, the value $v_{t}$ is retrieved through a query vector $q_{t+\tau}\in\mathbb{R}^{d}$. mLSTM uses a covariance update rule ($C_{t}=C_{t-1}+v_{t}\,k_{t}^{\top}$) to store the key-value pair. The covariance update rule is equivalent to the Fast Weight Programmer [72]. Later, a new variant has emerged [6]: a constant decay rate multiplied by $C_{t-1}$ and a constant learning rate multiplied by $v_{t}\,k_{t}^{\top}$. Similarly, in mLSTM, the forget gate corresponds to the decay rate, while the input gate corresponds to the learning rate. Furthermore, since the dot product between the query input and the normalizer state may approach zero, mLSTM uses the absolute value of the dot product and sets a lower bound to a threshold (e.g., 1). The cell state update rule for mLSTM is: